# Computing Chebyshev polynomials using the complex Remez algorithm

### v2

@@ 3.1.1 Regular polygon
then the set is invariant under rotations by an angle of $2\pi/m$ and hence Lemma 1 implies that
$T_{nm+l}^{\mathsf{E}}(z)=z^{l}Q_{n}^{\mathsf{E}}(z^{m}),\quad l=0,1,\dotsc,m-1,$ (18)
@@
$n=5$ 1.3090 1.2784 1.2135 1.5142
$n=10$ 1.1427 1.1298 1.1424 1.1736
$n=25$ 1.0549 1.0497 1.0554 1.0632
$n=50$ 1.0271 1.0245 1.0272 1.0314
$n=90$ 1.0135 1.0150 1.0173
$n=120$ 1.0104 1.0112 1.0130
Table 1: Widom factors corresponding to regular polygons, computed with an accuracy of $10^{-5}$ using Tang’s algorithm.
(a) $\mathcal{W}_{3n+l}(\mathsf{E}_{\Delta})$
(b) $\mathcal{W}_{4n+l}(\mathsf{E}_{\square})$
(c) $\mathcal{W}_{5n+l}(\mathsf{E}_{\pentagon})$
(d) $\mathcal{W}_{6n+l}(\mathsf{E}_{\hexagon})$
Figure 1: The different colors used in Figures 1(a)1(d) represent different values of $l\in\{0,1,\dotsc,m-1\}$ for $\mathcal{W}_{nm+l}$. The dotted blue line represents the value $\frac{m+2}{m}$ which relates to the Faber polynomials.
3.1.2 Hypocycloid
Examples of sets which are not quasi-circles are sets which contain outward pointing cusps on their boundary. With an “outward pointing cusp” we simply mean that the exterior angle at such a point is $2\pi$. For a pictorial representation the reader can consult Figures 5(g)-5(j) since examples of sets containing cusps are the $m$-cusped hypocycloids. These are the Jordan curves $\{\mathsf{H}_{m}\}$ defined via
$\mathsf{H}_{m}:=\left\{e^{i\theta}+\frac{e^{-i(m-1)\theta}}{m-1}:\theta\in[0,2\pi)\right\}.$ (19)
@@ 3.1.2 Hypocycloid
Clearly the sets $\mathsf{H}_{m}$ are invariant under rotations by $e^{2\pi i/m}$ and therefore Lemma 1 implies that
@@
$n=5$ 1.6959 1.5212 1.6445 2.48832
$n=10$ 1.4315 1.4078 1.4091 1.4744
$n=25$ 1.2518 1.2404 1.2493 1.2664
$n=50$ 1.1717 1.1626 1.1674 1.1759
$n=90$ 1.1113 1.1213 1.1269
Table 2: Widom factors corresponding to $m$-cusped Hypocycloids, computed with an accuracy of $10^{-5}$ using Tang’s algorithm.
(a) $\mathcal{W}_{3n+l}(\mathsf{H}_{3})$
(b) $\mathcal{W}_{4n+l}(\mathsf{H}_{4})$
(c) $\mathcal{W}_{5n+l}(\mathsf{H}_{5})$
(d) $\mathcal{W}_{6n+l}(\mathsf{H}_{6})$
Figure 2: The different colors used in Figures 2(a)-2(d) represent different values of $l\in\{0,1,\dotsc,m-1\}$ for $\mathcal{W}_{nm+l}$.
3.1.3 Circular Lunes
As a final example of the computation of Widom factors we consider the case of circular lunes, see Figures 5(e) and 5(f). Given $\alpha\in(0,2]$, we let
$\mathsf{C}_{\alpha}=\left\{\alpha\frac{1+\left(\frac{w-1}{w+1}\right)^{\alpha}}{1-\left(\frac{w-1}{w+1}\right)^{\alpha}}:|w|=1\right\}$ (20)
with vertices at ±$\pm\alpha$ and exterior angle $\alpha\pi$. The structure of such sets heavily depend on the value of the parameter $\alpha$. If $\alpha\in(0,1)$ then the set is non-convex while if $\alpha\in(1,2)$ the set is convex. The extreme cases are $\mathsf{C}_{1}=\mathbb{T}$ and $\mathsf{C}_{2}=[-2,2]$. We will consider two parameter values, namely $\alpha=\frac{1}{2}$ and $\alpha=\frac{3}{2}$ as they cover the cases of concavity and convexity. Irregardless of the parameter value of $\alpha$, the set $\mathsf{C}_{\alpha}$ is symmetric with respect to both axes. From Lemma 1 we conclude that
@@ 3.1.3 Circular Lunes
$n=10$ 1.0404 1.0619
$n=25$ 1.0341 1.0244
$n=30$ 1.0212 1.0203
$n=35$ 1.0261 1.0174
$n=45$ 1.0213 1.0135
Table 3: Widom factors corresponding to circular lunes, computed with an accuracy of $10^{-5}$ using Tang’s algorithm.
(a) $\mathcal{W}_{2n+l}(\mathsf{C}_{1/2})$
(b) $\mathcal{W}_{2n+l}(\mathsf{C}_{3/2})$
Figure 3: The different colors used in Figures 3(a) & 3(b) represent different values of $l\in\{0,1\}$ for $\mathcal{W}_{2n+l}$.
3.2 The Faber connection
Our initial interest in computing Chebyshev polynomials originated in studies of their zeros. One part of this study concerned Chebyshev polynomials on level curves corresponding to the exterior conformal map of a simply connected set $\mathsf{E}$. More precisely, if $\Phi$ is the exterior conformal map we investigated Chebyshev polynomials on the level curves
$\mathsf{E}^{r}:=\{z:|\Phi(z)|=r\}$
and found that the corresponding zeros of $T_{n}^{\mathsf{E}^{r}}$ seemed to converge for increasing $r$. By simultaneously plotting the zeros of the Faber polynomials, the picture became quite clear. The zeros of $T_{n}^{\mathsf{E}^{r}}$, as $r$ increased, appeared to accumulate at the zeros of the corresponding Faber polynomials. We investigate this possible relation numerically for lemniscates, hypocycloids and circular lunes.
3.2.1 Lemniscates
@@
$\mathsf{L}_{m}^{r}=\Big{\{}z:|z^{m}-1|=r^{m}\Big{\}}.$ (21)
A pictorial representation of such sets can be found in Figures 5(k) and 5(l). From (4) we gather that $\operatorname{Cap}(\mathsf{L}_{m}^{r})=r$ and since the polynomial $(z^{m}-1)^{n}$ saturates the lower bound in (3) we see that $T_{nm}^{\mathsf{L}_{m}^{r}}(z)=(z^{m}-1)^{n}$. For the remaining degrees we apply Lemma 1 to draw the conclusion that
$T_{nm+l}^{\mathsf{L}_{m}^{r}}(z)=z^{l}Q_{n}^{\mathsf{L}_{m}^{r}}(z^{m}),$ (22)
@@
  • If $r=1$, we write $\mathsf{L}_{m}^{1}=\mathsf{L}_{m}$ and in this case $\mathsf{L}_{m}$ is connected however its interior is not.
@@
Since $T_{mm}^{\mathsf{L}_{m}^{r}}(z)=(z^{m}-1)^{n}$, we see that $\mathcal{W}_{nm}(\mathsf{L}_{m}^{r})=1$ for any $n,m$ and $r$. The question is what the asymptotic behavior is for the remaining sequences of degrees. For $r>1$ it follows immediately from (8) that $\mathcal{W}_{n}(\mathsf{L}_{m}^{r})\rightarrow 1$ as $n\rightarrow\infty$ since the boundary is an analytic Jordan curve. If $10<r<$ 1 then it is known that $\limsup_{n\rightarrow\infty}\mathcal{W}_{n}(\mathsf{L}_{m}^{r})>1$, see [51]. The remaining case, when $r=1$, is handled by [5, Corollary 2] where it is shown that $\mathcal{W}_{n}(\mathsf{L}_{m}^{1})\rightarrow 1$ as $n\rightarrow\infty$.
@@
The conformal map taking ¯$\overline{\mathbb{C}}\setminus\mathsf{L}$ to $\{z:|z|>1\}$ with $\Phi(\infty)=\infty$ is given by
@@
$T_{3}^{\mathsf{L}^{r}}(z)=z\left(z^{2}-\frac{1}{5}\left(4-r^{4}+\sqrt{1+7r^{4}+r^{8}}\right)\right).$ (23)
@@
and hence we gather from (23) that $\lim_{r\rightarrow\infty}T_{3}^{\mathsf{L}^{r}}(z)=F_{3}^{\mathsf{L}}(z)$ uniformly on compact subsets of the complex plane. The question is whether this should be considered an anomaly or a potential link between Chebyshev polynomials and Faber polynomials. The natural procedure is of course to consider further examples. We do so numerically using Tang’s algorithm.
@@
and illustrate that this appears to tend to 00 with $r$. Such a difference is illustrated in Figure 4(a).
3.2.2 Hypocycloid
We continue the considerations concerning a possible relation between Faber polynomials and Chebyshev polynomials on level curves corresponding to conformal maps. We therefore return to the family of $m$-cusped hypocycloids $\{\mathsf{H}_{m}\}$. The Faber polynomials $F_{n}^{\mathsf{H}_{m}}$ can be computed using [24, Proposition 2.3].
@@ 3.2.2 Hypocycloid
$\mathsf{H}_{m}^{r}:=\left\{re^{i\theta}+\frac{(re^{i\theta})^{-(m-1)}}{m-1}:\theta\in[0,2\pi)\right\}.$
@@
as $r\rightarrow\infty$, we compute $\|F_{n}^{\mathsf{H}_{m}}-T_{n}^{\mathsf{H}_{m}^{r}}\|_{\infty}$ for $m=5$. The graph is illustrated in Figure 4(b).
3.2.3 Circular Lunes
We end the considerations of comparing Chebyshev polynomials to Faber polynomials by considering the case of circular lunes. As an example we consider the case where $\alpha=1/2$. In this case the canonical external conformal map $\Phi$ from the unbounded component of $\mathbb{C}\setminus\mathsf{C}_{1/2}$ to the exterior of the closed unit disk has the simple form
@@ 3.2.3 Circular Lunes
The computed difference $\|F_{n}^{\mathsf{C}_{1/2}}-T_{n}^{\mathsf{C}_{1/2}^{r}}\|_{\infty}$ for $n=11$ is illustrated in Figure 4(c).
(a) $r\mapsto\|F_{11}^{\mathsf{L}}-T_{11}^{\mathsf{L}^{r}}\|_{\infty}$
(b) $r\mapsto\|F_{11}^{\mathsf{H}_{5}}-T_{11}^{\mathsf{H}_{5}^{r}}\|_{\infty}$
(c) $r\mapsto\|T_{11}^{\mathsf{C}_{1/2}^{r}}-F_{11}^{\mathsf{C}_{1/2}}\|_{\infty}$
Figure 4: The figures represents $\|F_{n}^{\mathsf{E}}-T_{n}^{\mathsf{E}_{r}}\|_{\infty}$ as functions of $r>0$ for different $\mathsf{E}$ and $n=11$.
3.3 Zero distribution
Our final computations concerns computing the zeros of $T_{n}^{\mathsf{E}}$ for different compact sets $\mathsf{E}$. In Figures 5(a)-5(d) the zeros corresponding to $\mathsf{E}_{\Delta}$, $\mathsf{E}_{\square}$, $\mathsf{E}_{\pentagon}$ and $\mathsf{E}_{\hexagon}$ are computed. In Figures 5(e) and 5(f) the zeros of certain $T_{n}^{\mathsf{C}_{\alpha}}$ are illustrated for $\alpha=1/2$ and $\alpha=3/2$. In Figures 5(g)5(j) the zeros of certain $T_{n}^{\mathsf{H}_{m}}$ are computed for different values of $m$ and $n$. In Figures 5(k) and 5(l) the zeros corresponding to $T_{n}^{\mathsf{L}}$ and $T_{n}^{\mathsf{L}^{r}}$ are computed. To complement the plots of the zeros of $T_{n}^{\mathsf{L}^{r}}$ we also plot the zeros of Chebyshev polynomials corresponding to two different families of lemniscates. In particular lemniscates of the form
$\mathsf{E}_{4}^{r}:=\{z:|z^{4}-z^{2}|=r/4\},\quad\text{and}\quad\mathsf{E}_{3}^{r}:=\{z:|z^{3}+z+1|=r\}.$
The corresponding zero plots are given in Figures 5(m) and 5(n). We again stress the fact that the computations are approximative since we compute polynomials $P$ whose norms are close to $T_{n}^{\mathsf{E}}$.
(a) $T_{16}^{\mathsf{E}_{\Delta}}$, and $T_{62}^{\mathsf{E}_{\Delta}}$
(b) $T_{17}^{\mathsf{E}_{\square}}$, and $T_{82}^{\mathsf{E}_{\square}}$
(c) $T_{21}^{\mathsf{E}_{\pentagon}}$ and $T_{102}^{\mathsf{E}_{\pentagon}}$
(d) $T_{27}^{\mathsf{E}_{\hexagon}}$ and $T_{122}^{\mathsf{E}_{\hexagon}}$
(e) $T_{11}^{\mathsf{C}_{1/2}}$ and $T_{21}^{\mathsf{C}_{1/2}}$.
(f) $T_{10}^{\mathsf{C}_{3/2}}$, and $T_{50}^{\mathsf{C}_{3/2}}$
(g) $T_{31}^{\mathsf{H}_{3}}$, and $T_{47}^{\mathsf{H}_{3}}$
(h) $T_{41}^{\mathsf{H}_{4}}$, and $T_{62}^{\mathsf{H}_{4}}$
(i) $T_{51}^{\mathsf{H}_{5}}$ and $T_{77}^{\mathsf{H}_{5}}$
(j) $T_{61}^{\mathsf{H}_{6}}$ and $T_{92}^{\mathsf{H}_{6}}$
(k) $T_{11}^{\mathsf{L}}$ and $T_{41}^{\mathsf{L}}$.
(l) $T_{11}^{\mathsf{L}^{2}}$ and $T_{41}^{\mathsf{L}^{2}}$. Dotted: $\mathsf{L}$.
(m) $T_{21}^{\mathsf{E}_{4}^{5/4}}$ and $T_{37}^{\mathsf{E}_{4}^{5/4}}$. Dotted: $\mathsf{E}_{4}^{1/4}$.
(n) $T_{25}^{\mathsf{E}_{3}^{2}}$ and $T_{40}^{\mathsf{E}_{3}^{2}}$. Dotted: $\mathsf{E}_{3}^{\sqrt{31/27}}$.
Figure 5: Zeros of Chebyshev polynomials corresponding to different sets. In Figures 5(l), 5(m) and 5(n) the Chebyshev polynomials are associated with the filled in curve. The dotted curve corresponds to the critical level curve.
4 Discussion

### v3

@@ 3.1.1. Regular polygons
then the set is invariant under rotations by an angle of $2\pi/m$ and hence Lemma 3.1 implies that
(21) $T_{nm+l}^{\mathsf{E}}(z)=z^{l}Q_{n}^{\mathsf{E}}(z^{m}),\quad l=0,1,\dotsc,m-1,$
@@
$n=5$ 1.30901051 1.27841716 1.21350890 1.51420435
$n=10$ 1.14268975 1.12981144 1.14236706 1.17363458
$n=25$ 1.05488942 1.04969579 1.05544736 1.06322465
$n=50$ 1.02708221 1.02449420 1.02724022 1.03142381
$n=90$ 1.01495704 1.01352749 1.01502916 1.01733310
$n=120$ 1.01119706 1.01012748 1.01124879 1.01297657
Table 1. Widom factors corresponding to regular polygons, computed with an accuracy of $10^{-10}$ using Tang’s algorithm.
(a) Widom factors of an equilateral triangle.
(b) Widom factors of a square.
(c) Widom factors of a regular pentagon.
(d) Widom factors of a regular hexagon.
Figure 1. The difference between $\mathcal{W}_{n}$ and $1$ for different $m$-gons. The dotted blue line represents the value $\frac{m+2}{m}-1$ which relates to the Faber polynomials. The Tang accuracy is $10^{-7}$.
3.1.2. Hypocycloids
Examples of sets which are not quasi-circles are sets which contain cusps on their boundary. With an “outward pointing cusp” we simply mean that the exterior angle at such a point is $2\pi$. For a pictorial representation the reader can consult Figures 7(a)-7(d) illustrating $m$-cusped hypocycloids. These are the Jordan curves $\{\mathsf{H}_{m}\}$ defined via
(22) $\mathsf{H}_{m}\coloneqq\left\{e^{i\theta}+\frac{e^{-i(m-1)\theta}}{m-1}:\theta\in[0,2\pi)\right\}.$
@@ 3.1.2. Hypocycloids
The sets $\mathsf{H}_{m}$ are clearly invariant under rotations by $e^{2\pi i/m}$ and therefore Lemma 3.1 implies that
@@
$n=5$ 1.69594045 1.52124467 1.64453125 2.48832000
$n=10$ 1.43149779 1.40782752 1.40910966 1.47443526
$n=25$ 1.25181361 1.24037027 1.24939099 1.26640101
$n=50$ 1.17181774 1.16257852 1.16744376 1.17588712
$n=90$ 1.12543411 1.11836230 1.12140197 1.12692589
$n=120$ 1.10776251 1.10161806 1.10410341 1.10869246
Table 2. Widom factors corresponding to $m$-cusped Hypocycloids, computed with an accuracy of $10^{-10}$ using Tang’s algorithm.
(a) Widom factors of $\mathsf{H}_{3}$.
(b) Widom factors of $\mathsf{H}_{4}$.
(c) Widom factors of $\mathsf{H}_{5}$.
(d) Widom factors of $\mathsf{H}_{6}$.
Figure 2. The difference between $\mathcal{W}_{n}$ and $1$ on different $m$-cusped Hypocycloids. The Tang accuracy is $10^{-7}.$
3.1.3. Circular lunes
As a final example of the computation of Widom factors we consider the case of circular lunes, see Figures 6(a) and 6(b). Given $\alpha\in(0,2]$, we let
(23) $\mathsf{C}_{\alpha}\coloneqq\left\{\alpha\frac{1+\left(\frac{w-1}{w+1}\right)^{\alpha}}{1-\left(\frac{w-1}{w+1}\right)^{\alpha}}:|w|=1\right\}$
with vertices at ±$\pm\alpha$ and exterior angle $\alpha\pi$. Note that $\operatorname{Cap}(\mathsf{C}_{\alpha})=1$. The structure of such sets depends on the value of the parameter $\alpha$. $\mathsf{C}_{\alpha}$ is non-convex if $\alpha\in(0,1)$ and convex if $\alpha\in(1,2)$. The extreme cases are $\mathsf{C}_{1}=\mathbb{T}$ and $\mathsf{C}_{2}=[-2,2]$. We will consider two parameter values, namely $\alpha=\frac{1}{2}$ and $\alpha=\frac{3}{2}$ as they have different convexity properties. Regardless of the parameter value of $\alpha$, the set $\mathsf{C}_{\alpha}$ is symmetric with respect to both axes. From Lemma 3.1 we conclude that
@@ 3.1.3. Circular lunes
$n=5$ 1.10286958 1.12569879
$n=10$ 1.03696888 1.06185388
$n=25$ 1.03405451 1.02444481
$n=50$ 1.01442556 1.01215983
$n=90$ 1.00936347 1.00673877
$n=120$ 1.00749065 1.00505004
Table 3. Widom factors corresponding to circular lunes, computed with an accuracy of $10^{-10}$ using Tang’s algorithm.
(a) Widom factors of $\mathsf{C}_{1/2}$.
(b) Widom factors of $\mathsf{C}_{3/2}$.
Figure 3. The difference betweeen $\mathcal{W}_{n}$ and $1$ on two different circular lunes. The Tang accuracy is $10^{-7}.$
3.2. The Faber connection
Our initial motivation for computing Chebyshev polynomials originated in obtaining a better understanding of the behavior of their zeros. One case that we wished to better understand concerned Chebyshev polynomials on equipotential curves. More precisely, if $\Phi$ is the exterior conformal map of a connected set $\mathsf{E}$, we investigated Chebyshev polynomials on the level curves
$\mathsf{E}^{r}\coloneqq\{z:|\Phi(z)|=r\}$
and found that the corresponding zeros of $T_{n}^{\mathsf{E}^{r}}$ seemed to converge as $r$ increases. By simultaneously plotting the zeros of the Faber polynomials, the picture became quite clear. The zeros of $T_{n}^{\mathsf{E}^{r}}$, as $r$ increased, appeared to converge toward the zeros of the corresponding Faber polynomials. We now investigate this possible relation numerically for lemniscates, hypocycloids and circular lunes.
3.2.1. Lemniscates
@@
(24) $\mathsf{L}_{m}^{r}\coloneqq\Big{\{}z:|z^{m}-1|=r^{m}\Big{\}}.$
A pictorial representation of such sets can be found in Figures 8(a) and 8(b). From (4) we observe that $\operatorname{Cap}(\mathsf{L}_{m}^{r})=r$ and since the polynomial $(z^{m}-1)^{n}$ saturates the lower bound in (3) we see that $T_{nm}^{\mathsf{L}_{m}^{r}}(z)=(z^{m}-1)^{n}$. For the remaining degrees we apply Lemma 3.1 to conclude
(25) $T_{nm+l}^{\mathsf{L}_{m}^{r}}(z)=z^{l}Q_{n}^{\mathsf{L}_{m}^{r}}(z^{m}),$
@@
  • If $r=1$, we write $\mathsf{L}_{m}^{1}=\mathsf{L}_{m}$ and in this case $\mathsf{L}_{m}$ is connected, but its interior is not.
@@
Since $T_{mm}^{\mathsf{L}_{m}^{r}}(z)=(z^{m}-1)^{n}$, we see that $\mathcal{W}_{nm}(\mathsf{L}_{m}^{r})=1$ for any $n,m$ and $r$. The question is what the asymptotic behavior is for the remaining sequences of degrees. For $r>1$ it follows immediately from (8) that $\mathcal{W}_{n}(\mathsf{L}_{m}^{r})\rightarrow 1$, as $n\rightarrow\infty$ since the boundary is an analytic Jordan curve. If $10<r<$ 1 then it is known that $\limsup_{n\rightarrow\infty}\mathcal{W}_{n}(\mathsf{L}_{m}^{r})>1$, see [54, Theorem 2]. The remaining case, when $r=1$, is handled by [6, Corollary 2] where it is shown that $\mathcal{W}_{n}(\mathsf{L}_{m}^{1})\rightarrow 1$ as $n\rightarrow\infty$.
@@
The conformal mapping from ¯$\overline{\mathbb{C}}\setminus\mathsf{L}$ to $\{z:|z|>1\}$ with $\Phi(\infty)=\infty$ is given by
@@
(26) $T_{3}^{\mathsf{L}^{r}}(z)=z\left(z^{2}-\frac{1}{5}\left(4-r^{4}+\sqrt{1+7r^{4}+r^{8}}\right)\right).$
@@
and hence we gather from (26) that $T_{3}^{\mathsf{L}^{r}}(z)\rightarrow F_{3}^{\mathsf{L}}(z)$ as $r\rightarrow\infty$, uniformly on compact subsets of the complex plane. The question is whether this should be considered an anomaly or a potential link between Chebyshev polynomials and Faber polynomials. The natural procedure is of course to consider further examples. We do so numerically using Tang’s algorithm.
@@
and illustrate that this appears to tend to 00 with $r\rightarrow\infty$, see Figure 4.
3.2.2. Hypocycloids
We continue the considerations concerning a possible relation between Faber polynomials and Chebyshev polynomials on level curves corresponding to conformal maps. We therefore return to the family of $m$-cusped hypocycloids $\{\mathsf{H}_{m}\}$. The Faber polynomials $F_{n}^{\mathsf{H}_{m}}$ can be computed using [26, Proposition 2.3].
@@ 3.2.2. Hypocycloids
$\mathsf{H}_{m}^{r}\coloneqq\left\{re^{i\theta}+\frac{(re^{i\theta})^{-(m-1)}}{m-1}:\theta\in[0,2\pi)\right\}.$
@@
as $r\rightarrow\infty$, we compute $\|F_{n}^{\mathsf{H}_{m}}-T_{n}^{\mathsf{H}_{m}^{r}}\|_{\infty}$ for $m=5$, and illustrate it in Figure 4.
3.2.3. Circular lunes
We end the considerations of comparing Chebyshev polynomials to Faber polynomials by considering the case of circular lunes. As an example, we consider the case where $\alpha=1/2$. In this case the canonical external conformal map $\Phi$ from the unbounded component of $\mathbb{C}\setminus\mathsf{C}_{1/2}$ to the exterior of the closed unit disk has the simple form
@@ 3.2.3. Circular lunes
The computed difference $\|F_{n}^{\mathsf{C}_{1/2}}-T_{n}^{\mathsf{C}_{1/2}^{r}}\|_{\infty}$ for $n=11$ is illustrated in Figure 4.
Figure 4. The figures illustrate $\|F_{n}^{\mathsf{E}}-T_{n}^{\mathsf{E}_{r}}\|_{\infty}$ as functions of $r>1$ for different $\mathsf{E}$ and $n=11$ using Tang accuracy of $10^{-40}$. For $\mathsf{L}^{r}$ and $\mathsf{C}_{1/2}^{r}$ the estimated convergence rate is around $O(r^{-4})$. For $\mathsf{H}_{5}^{r}$ the estimated convergence rate is $O(r^{-10})$.
3.3. Zero distribution
Our final computations concern zeros of $T_{n}^{\mathsf{E}}$ for different compact sets $\mathsf{E}$. In Figures 5(a)-5(d) the zeros corresponding to $\mathsf{E}_{\Delta}$, $\mathsf{E}_{\square}$, $\mathsf{E}_{\pentagon}$ and $\mathsf{E}_{\hexagon}$ are displayed. In Figures 6(a) and 6(b) the zeros of $T_{n}^{\mathsf{C}_{\alpha}}$ are illustrated for $\alpha=1/2$ and $\alpha=3/2$. In Figures 7(a)7(d) the zeros of $T_{n}^{\mathsf{H}_{m}}$ are illustrated for different values of $m$ and $n$. In Figures 8(a) and 8(b) the zeros corresponding to $T_{n}^{\mathsf{L}}$ and $T_{n}^{\mathsf{L}^{r}}$ are illustrated. To complement the plots of the zeros of $T_{n}^{\mathsf{L}^{r}}$ we also plot the zeros of Chebyshev polynomials corresponding to
$\mathsf{A}^{r}\coloneqq\{z:|z^{3}+z+1|=r\},\quad\text{and}\quad\mathsf{B}^{r}\coloneqq\{z:|z^{4}-z^{2}|=r\}.$
We introduce the constants
(28) $r_{3}=\sqrt{31/27},\quad\text{ and }\quad r_{4}=1/4$
since $\mathsf{A}^{r_{3}}$ and $\mathsf{B}^{r_{4}}$ correspond to the critical case where the lemniscates have self-intersections. The corresponding zero plots are given in Figures 8(c) and 8(d).
(a) $T_{16}^{\mathsf{E}_{\Delta}}$ and $T_{62}^{\mathsf{E}_{\Delta}}$.
(b) $T_{17}^{\mathsf{E}_{\square}}$ and $T_{82}^{\mathsf{E}_{\square}}$.
(c) $T_{21}^{\mathsf{E}_{\pentagon}}$ and $T_{102}^{\mathsf{E}_{\pentagon}}$.
(d) $T_{27}^{\mathsf{E}_{\hexagon}}$ and $T_{122}^{\mathsf{E}_{\hexagon}}$.
Figure 5. Zeros of Chebyshev polynomials corresponding to $m$-gons.
(a) $T_{11}^{\mathsf{C}_{1/2}}$ and $T_{21}^{\mathsf{C}_{1/2}}$.
(b) $T_{10}^{\mathsf{C}_{3/2}}$ and $T_{50}^{\mathsf{C}_{3/2}}$.
Figure 6. Zeros of Chebyshev polynomials corresponding to circular lunes.
(a) $T_{31}^{\mathsf{H}_{3}}$ and $T_{47}^{\mathsf{H}_{3}}$.
(b) $T_{41}^{\mathsf{H}_{4}}$ and $T_{62}^{\mathsf{H}_{4}}$.
(c) $T_{51}^{\mathsf{H}_{5}}$ and $T_{77}^{\mathsf{H}_{5}}$.
(d) $T_{61}^{\mathsf{H}_{6}}$ and $T_{92}^{\mathsf{H}_{6}}$.
Figure 7. Zeros of Chebyshev polynomials corresponding to $m$-cusped Hypocycloids.
(a) $T_{11}^{\mathsf{L}_{2}}$ and $T_{41}^{\mathsf{L}_{2}}$.
(b) $T_{11}^{\mathsf{L}_{2}^{2}}$ and $T_{41}^{\mathsf{L}_{2}^{2}}$. Dotted: $\mathsf{L}_{2}$.
(c) $T_{21}^{\mathsf{B}^{5/4}}$ and $T_{37}^{\mathsf{B}^{5/4}}$. Dotted: $\mathsf{B}^{r_{4}}$.
(d) $T_{25}^{\mathsf{A}^{2}}$ and $T_{40}^{\mathsf{A}^{2}}$. Dotted: $\mathsf{A}^{r_{3}}$.
Figure 8. Zeros of Chebyshev polynomials corresponding to lemniscates.
4. Discussion